OpenStudy (anonymous):

simplify the square root of 32

OpenStudy (anonymous):

simplify\[\sqrt{32}\]

OpenStudy (anonymous):

do you want the actual simplified number?

OpenStudy (anonymous):

the whole number is 5.65685425

OpenStudy (anonymous):

you have either 5.7, 5.66

OpenStudy (anonymous):

well like \[x \sqrt{y}\]

OpenStudy (anonymous):

breaking down the square into two numbers

OpenStudy (anonymous):

in that case, hang on a sec

OpenStudy (anonymous):

8\[\sqrt{4}\]

OpenStudy (anonymous):

eight times the square root of 2

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

no problem